On Prior Navigation Knowledge in Multi Sensor Indoor Localisation

Frank Ebner, Toni Fetzer and Frank Deinzer

Faculty of Computer Science and Business Information Systems University of Applied Sciences Würzburg-Schweinfurt Würzburg, Germany {frank.ebner, toni.fetzer, frank.deinzer}@fhws.de

Marcin Grzegorzek

Pattern Recognition Group University of Siegen Siegen, Germany {marcin.grzegorzek}@uni-siegen.de

Abstract—

I. Introduction

Smoothing Smoothing Smoothing

II. Related Work III. Recursive State Estimation IV. Transition V. Smoothing

Test

VI. Experiments

ddd [1] dddd

VII. Conclusion References

[1] V. Honkavirta, T. Perl, S. Ali-Lytty, and R. Piche, “A comparative survey of wlan location fingerprinting methods,” Workshop on Positioning, Navigation and Communication, vol. 6, no. 1, pp. 243–251, 2009.